2016 FICA Tax Rates1
Employee's Rates Matching Rates Paid by
the Employer Self-Employed Rates
7.65% on first $118,500 of income
1.45% on income above $118,500
7.65% on first $118,500 paid in wages
1.45% on wages above $118,500
13.3% on first $118,500 of net earnings
2.9% on earnings above $118,500
1FICA taxes include Social Security and Medicare. The Social Security tax applies to the first $118,500 of income, while the Medicare tax applies to all income.

Suppose Jamison has $155,100 of income from work and is not self-employed. How much will Jamison have to pay in FICA taxes?

Answer by Theo(13342)

if you file as self employed, you are paying the employer tax rate plus the employee tax rate.

that winds up being double the employee tax rate.

it doesn't figure into this problem, but i just wanted to let you know that your 13.3% for social security and medicare up to 118,500 should be shown as 15.3%, not 13.3%.

here's the 2016 fica tax rates from the internet.

https://sfcontroller.org/sites/default/files/Documents/payroll/EmploymentTaxWage2016.pdf

your problem states that Suppose Jamison has $155,100 of income from work and is not self-employed. How much will Jamison have to pay in FICA taxes?

his income is 155,100.

he will pay 7.65% on the first 118,500 and 1.45% on the rest.

.0765 * 118,500 = 9,065.25.

the rest is 155,100 minus 118,500 = 36,600.

he pays .0145 * 36,600 = 530.7 on that.

the total that he has to pay is 9,065.25 + 530.7 = 9,595.95.